# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 1.3 Contribution
  • We propose a novel SLRQA model to address image processing problems with noise by maximizing the a posteriori probability estimate. The proposed SLRQA takes advantage of quaternion representation, low-rankness, and sparsity without requiring an estimation of the rank.
  • We give a proximal linearized alternating direction method of multipliers (PL-ADMM) algorithm. Under mild conditions, we prove that the sequence generated by the algorithm converges to a stationary point. To the best of our knowledge, this is the first convergence analysis result of an algorithm for a quaternion-based image processing model.
  • For noise-free observations, we propose a variant of SLRQA called SLRQA-NF. Inspired by PL-ADMM, an algorithm named PL-ADMM-NF is presented to solve the SLRQA-NF model.
@@ 1.4 Organization
The rest of this paper is organized as follows. In Section 2, the preliminaries and notations are introduced. The PL-ADMM algorithm framework for solving SLRQA model is given in Section 3. The global convergence is established in Section 4. In Section 5, the PL-ADMM-NF algorithm for solving SLRQA-NF model is presented. In Section 6, extensive numerical experiments are conducted to demonstrate the performance of the models and algorithms by comparing them with other state-of-the-art methods. Finally, the conclusion is given in Section 7.
@@ 2 Preliminaries
Throughout this paper, scalars, real vectors, real matrices, and real tensor are respectively denoted by lowercase letters $x$, boldface lowercase letters $\mathbf{x}$, boldface capital letters $\mathbf{X}$, and boldface capital letters with a tilde $\widetilde{\mathbf{X}}$ respectively. $\mathbb{H}$ denotes the quaternion algebra proposed in [16]. A dot above a variable $\dot{q}$ represents a quaternion number. A quaternion number consists of one real part and three imaginary parts:
$\dot{q}=x_{s}+x_{i}\mathbf{i}+x_{j}\mathbf{j}+x_{k}\mathbf{k}.$
@@
where $x_{s},x_{i},x_{j},x_{k}\in\mathbb{R}$, and $\mathbf{i},\mathbf{j},\mathbf{k}$ are three imaginary units. The quaternion $\dot{q}$ is called a pure quaternion if and only if it has a zero real part, $\mathbf{\;i.e.\;}x_{s}=0$. The real part of $\dot{q}$ is $\text{Re}(\dot{q})=x_{s}$. The conjugate $\overline{\dot{q}}$ is defined as $\overline{\dot{q}}=x_{s}-x_{i}\mathbf{i}-x_{j}\mathbf{j}-x_{k}\mathbf{k}$. The modulus $|\dot{q}|$ is defined as $|\dot{q}|=\sqrt{x_{s}^{2}+x_{i}^{2}+x_{j}^{2}+x_{k}^{2}}$. The inverse of $\dot{q}$ is given by $\dot{q}^{-1}=\dfrac{\overline{\dot{q}}}{|\dot{q}|^{2}}$. For a quaternion matrix ${\mathbf{\dot{X}}},$ ${\mathbf{\dot{X}}}_{ij}$ denotes the $i$-th row, $j$-th coloum element of ${\mathbf{\dot{X}}}_{ij}$ and ${\mathbf{\dot{X}}}_{ijk}$ to denote the $i$-th row, $j$-th and $k$-th imaginary units. The Frobenius norm of ${\mathbf{\dot{X}}}$ is defined as $\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}|{\mathbf{\dot{X}}}_{ij}|^{2}}$. Accordingly, the conjugate operator $\overline{\mathbf{\dot{A}}}$ , transpose operator $\mathbf{\dot{A}}^{\mathrm{T}}$ and the conjugate transpose operator $\mathbf{\dot{A}}^{*}$ are defined as $\overline{\mathbf{\dot{A}}}=(\overline{\dot{a}})_{ij},\mathbf{\dot{A}}^{\mathrm{T}}=(\dot{a})_{ji}$ and $\mathbf{\dot{A}}^{*}=(\overline{\dot{a}})_{ji}$, respectively. Furthermore, the multiplication rule of quaternion is given by: $\mathbf{i}\cdot 1=\mathbf{i},\quad\mathbf{j}\cdot 1=\mathbf{j},\quad\mathbf{k}\cdot 1=\mathbf{k}.\quad$ $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-1,\mathbf{i}\mathbf{j}=-\mathbf{j}\mathbf{i}=\mathbf{k},\,\mathbf{j}\mathbf{k}=-\mathbf{k}\mathbf{j}=\mathbf{i},\,\mathbf{k}\mathbf{i}=-\mathbf{i}\mathbf{k}=\mathbf{j}$. Note that the multiplication of two quaternions $\dot{q}_{1}$ and $\dot{q}_{2}$ is not commutative, i.e., $\dot{q}_{1}\dot{q}_{2}\neq\dot{q}_{2}\dot{q}_{1}$ in general. $\mathbf{\dot{Q}}$ is called a unitary matrix if and only if $\mathbf{\dot{Q}}\mathbf{\dot{Q}}^{*}=\mathbf{\dot{Q}}^{*}\mathbf{\dot{Q}}=\mathbf{I},$ where $\mathbf{I}$ denotes a real identity matrix. The inner product of two quaternion matrices $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$ is defined as $\left\langle\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\right\rangle:=\text{tr}(\mathbf{\dot{G}}_{1}^{*}\mathbf{\dot{G}}_{2})$. The image of a given matrix $\mathbf{\dot{G}}_{1}$ is denoted by ${\rm range}(\mathbf{\dot{G}}_{1})$. Given an operator $\mathcal{W}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$, the weighted norm is defined as $\|{\mathbf{\dot{X}}}\|_{\mathcal{W}}:=\left\langle\mathcal{W}({\mathbf{\dot{X}}}),{\mathbf{\dot{X}}}\right\rangle$. The operator norm is defined as $\|\mathcal{W}\|:=\sup_{{\mathbf{\dot{X}}}}\frac{\|\mathcal{W}({\mathbf{\dot{X}}})\|_{\mathrm{F}}}{\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}}$. Let $\mathcal{W}^{\#}$ denote the adjoint operator of $\mathcal{W}$, i.e., $\left\langle\mathcal{W}(\mathbf{\dot{G}}_{1}),\mathbf{\dot{G}}_{2}\right\rangle=\left\langle\mathbf{\dot{G}}_{1},\mathcal{W}^{\#}(\mathbf{\dot{G}}_{2})\right\rangle$ holds for any $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$. If $\mathcal{W}$ is given by a quaternion matrix, then $\mathcal{W}^{\#}$ is the conjugate transpose of the matrix. For $\widetilde{\mathbf{\dot{G}}}:=(\mathbf{\dot{G}}_{1},\cdots,\mathbf{\dot{G}}_{p})\in\mathbb{H}^{n_{1}\times m_{1}}\times\cdots\mathbb{H}^{n_{p}\times m_{p}},$ $|||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}^{2}}$. Hence we have $\sum_{i=1}^{p}\frac{1}{\sqrt{p}}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}\leq\||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}\leq\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}},$ where the first equality follows from Cauchy-Schwarz inequality.
@@
3 An Algorithm for SLRQA model
@@
1:A noisy observed image $\mathbf{\dot{Y}}$; an initial iterate ${\mathbf{\dot{X}}}_{0}:=\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; two sequences of regularization parameters $\{L_{1,k}\}$ and $\{L_{2,k}\}$ satisfying $\sup_{k\geq 0}{L_{i,k}}<\infty$ and $\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$;
2:A recovered image ${\mathbf{\dot{X}}}_{*}$;
@@
${\mathbf{\dot{X}}}_{k+1}\in$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\;\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$
$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}_{k}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2};$
@@
$\mathbf{\dot{W}}_{k+1}=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}$ $\;\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{W}}-(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ $\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}$
@@
In Step 5, ${\mathbf{\dot{X}}}_{k+1}$ is obtained by minimizing $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed $\mathbf{\dot{W}}_{k}$ and $\mathbf{\dot{\Lambda}}_{k}$. This subproblem does not yield a closed-form solution in general and therefore may be solved by nonsmooth algorithms such as a subgradient method or the linearized Bregman algorithm in [4]. If the linear operator $\mathcal{A}$ in (2) is the identity and all pixels of the noisy image $\mathbf{\dot{Y}}$ are observed, i.e., $\Omega$ indicates the whole image, then the ${\mathbf{\dot{X}}}$-subproblem can be solved efficiently. Specifically, the subproblem can be reformulated as
${\mathbf{\dot{X}}}_{k+1}=$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ (10)
$\qquad\qquad\frac{\tau^{2}(\beta+L_{1,k})+1}{2\tau^{2}}\left\|{\mathbf{\dot{X}}}-\frac{\tau^{2}}{\tau^{2}(\beta+L_{1,k})+1}\left(\frac{1}{\tau^{2}}\mathbf{\dot{Y}}+\beta\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k}-\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})+L_{1,k}{\mathbf{\dot{X}}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
If $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sigma_{i}({\mathbf{\dot{X}}}),$ the ${\mathbf{\dot{X}}}$-subproblem has a closed-form solution by using the soft thresholding method on the singular values [4]. When a nonconvex surrogate function $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ is used, the subproblem of ${\mathbf{\dot{X}}}$ can be solved via a convex-concave procedure(CCP) in [20]. For completeness, CCP is stated in Algorithm 2. Lemma 3.1 is crucial in CCP and it transforms the matrix optimization problem of ${\mathbf{\dot{X}}}$ into an optimization problem of singular values.
Lemma 3.1 ([10]).
@@
$\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\,\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}$ (11)
@@
$\bm{\sigma_{*}}=\mathop{\mathrm{arg\,min}}_{\bm{\sigma}\,\geq 0}\phi(\bm{\sigma},\gamma)+\frac{\mu}{2}\|\bm{\sigma}-\bm{\sigma_{\bm{\hat{{\mathbf{\dot{X}}}}}}}\|_{2}^{2}.$ (12)
In the $k$-th iteration of Algorithm 2, the function $\phi(\bm{\sigma},\gamma)$ is approximated by a linearized function $\phi(\bm{\sigma}_{k},\gamma)+\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)^{\mathrm{T}}\bm{\sigma}$. It follows that the update of $\bm{\sigma}$ in Problem (12) at $\bm{\sigma}_{k}$ is given by
$\bm{\sigma}_{k+1}=\max\left\{\bm{\sigma}_{\bm{\hat{{\mathbf{\dot{X}}}}}}-\frac{\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)}{\mu},\bm{0}\right\}.$ (13)
It has been shown in [20, Theorem 4] that any limit points of $\{\bm{\sigma}_{k}\}_{k=1}^{\infty}$ are stationary points of (12).

### v2

@@ 1.3 Contribution
  • We propose a novel model called SLRQA to address image processing problems with noise by maximizing the a posteriori probability estimate. Unlike the existing image processing models, the proposed SLRQA (2) is a quaternion model combining low rank and sparsity, providing flexibility to choose surrogate functions and not needing an initial rank estimate. We propose a PL-ADMM algorithm to solve SLRQA and its global convergence is guaranteed.
  • When the observations are noise-free, we propose a model named SLRQA-NF (4) and a corresponding algorithm named PL-ADMM-NF. The global convergence of PL-ADMM-NF is established under a newly proposed assumption. To the best of our knowledge, this is the first nonconvex ADMM-type algorithm without the “range assumption” in (12) and still guarantees global convergence.
@@ 1.4 Organization
The rest of this paper is organized as follows. In Section 2, the preliminaries and notations are introduced. The PL-ADMM algorithm for solving SLRQA and PL-ADMM-NF algorithm for solving SLRQA-NF are presented in Section 3. The global convergence of these two algorithms are established in Section 4. In Section 5, extensive numerical experiments are conducted to demonstrate the performance of the models and algorithms by comparing them with other state-of-the-art methods. Finally, the conclusion is given in Section 6.
@@ 2 Preliminaries
Throughout this paper, scalars, real vectors, real matrices, and real tensors are respectively denoted by lowercase letters $x$, boldface lowercase letters $\mathbf{x}$, boldface capital letters $\mathbf{X}$, and boldface capital letters with a tilde $\widetilde{\mathbf{X}}$ respectively. $\mathbb{H}$ denotes the quaternion algebra proposed in [17]. A dot above a variable $\dot{q}$ represents a quaternion number. A quaternion number consists of one real part and three imaginary parts: $\dot{q}=x_{s}+x_{i}\mathbf{i}+x_{j}\mathbf{j}+x_{k}\mathbf{k}.$ where $x_{s},x_{i},x_{j},x_{k}\in\mathbb{R}$, and $\mathbf{i},\mathbf{j},\mathbf{k}$ are three imaginary units. The quaternion $\dot{q}$ is called a pure quaternion if and only if it has a zero real part, $\mathbf{\;i.e.\;}x_{s}=0$. The real part of $\dot{q}$ is $\text{Re}(\dot{q})=x_{s}$. The conjugate $\overline{\dot{q}}$ is defined as $\overline{\dot{q}}=x_{s}-x_{i}\mathbf{i}-x_{j}\mathbf{j}-x_{k}\mathbf{k}$. The modulus $|\dot{q}|$ is defined as $|\dot{q}|=\sqrt{x_{s}^{2}+x_{i}^{2}+x_{j}^{2}+x_{k}^{2}}$. The inverse of $\dot{q}$ is given by $\dot{q}^{-1}=\dfrac{\overline{\dot{q}}}{|\dot{q}|^{2}}$. For a quaternion matrix ${\mathbf{\dot{X}}},$ ${\mathbf{\dot{X}}}_{ij}$ denotes the $i$-th row, $j$-th coloum element of ${\mathbf{\dot{X}}}_{ij}$ and ${\mathbf{\dot{X}}}_{ijk}$ to denote the $i$-th row, $j$-th and $k$-th imaginary units. The Frobenius norm of ${\mathbf{\dot{X}}}$ is defined as $\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}|{\mathbf{\dot{X}}}_{ij}|^{2}}$. Accordingly, the conjugate operator $\overline{\mathbf{\dot{A}}}$ , transpose operator $\mathbf{\dot{A}}^{\mathrm{T}}$ and the conjugate transpose operator $\mathbf{\dot{A}}^{*}$ are defined as $\overline{\mathbf{\dot{A}}}=(\overline{\dot{a}})_{ij},\mathbf{\dot{A}}^{\mathrm{T}}=(\dot{a})_{ji}$ and $\mathbf{\dot{A}}^{*}=(\overline{\dot{a}})_{ji}$, respectively. Furthermore, the multiplication rule of quaternion is given by: $\mathbf{i}\cdot 1=\mathbf{i},\quad\mathbf{j}\cdot 1=\mathbf{j},\quad\mathbf{k}\cdot 1=\mathbf{k}.\quad$ $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-1,\mathbf{i}\mathbf{j}=-\mathbf{j}\mathbf{i}=\mathbf{k},\,\mathbf{j}\mathbf{k}=-\mathbf{k}\mathbf{j}=\mathbf{i},\,\mathbf{k}\mathbf{i}=-\mathbf{i}\mathbf{k}=\mathbf{j}$. Note that the multiplication of two quaternions $\dot{q}_{1}$ and $\dot{q}_{2}$ is not commutative, i.e., $\dot{q}_{1}\dot{q}_{2}\neq\dot{q}_{2}\dot{q}_{1}$ in general. $\mathbf{\dot{Q}}$ is called a unitary matrix if and only if $\mathbf{\dot{Q}}\mathbf{\dot{Q}}^{*}=\mathbf{\dot{Q}}^{*}\mathbf{\dot{Q}}=\mathbf{I},$ where $\mathbf{I}$ denotes a real identity matrix. The inner product of two quaternion matrices $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$ is defined as $\left\langle\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\right\rangle:=\text{tr}(\mathbf{\dot{G}}_{1}^{*}\mathbf{\dot{G}}_{2})$. The image of a given matrix $\mathbf{\dot{G}}_{1}$ is denoted by ${\rm range}(\mathbf{\dot{G}}_{1})$. Given an operator $\mathcal{W}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$, the weighted norm is defined as $\|{\mathbf{\dot{X}}}\|_{\mathcal{W}}:=\left\langle\mathcal{W}({\mathbf{\dot{X}}}),{\mathbf{\dot{X}}}\right\rangle$. The operator norm is defined as $\|\mathcal{W}\|:=\sup_{{\mathbf{\dot{X}}}}\frac{\|\mathcal{W}({\mathbf{\dot{X}}})\|_{\mathrm{F}}}{\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}}$. Let $\mathcal{W}^{\#}$ denote the adjoint operator of $\mathcal{W}$, i.e., $\left\langle\mathcal{W}(\mathbf{\dot{G}}_{1}),\mathbf{\dot{G}}_{2}\right\rangle=\left\langle\mathbf{\dot{G}}_{1},\mathcal{W}^{\#}(\mathbf{\dot{G}}_{2})\right\rangle$ holds for any $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$. If $\mathcal{W}$ is given by a quaternion matrix, then $\mathcal{W}^{\#}$ is the conjugate transpose of the matrix. For $\widetilde{\mathbf{\dot{G}}}:=(\mathbf{\dot{G}}_{1},\cdots,\mathbf{\dot{G}}_{p})\in\mathbb{H}^{n_{1}\times m_{1}}\times\cdots\mathbb{H}^{n_{p}\times m_{p}},$ $|||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}^{2}}$. Hence we have $\sum_{i=1}^{p}\frac{1}{\sqrt{p}}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}\leq\||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}\leq\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}},$ where the first equality follows from the Cauchy-Schwarz inequality.
@@
3 Algorithm framework
In this section, two algorithms, i.e. PL-ADMM and PL-ADMM-NF, are proposed to solve Model (2) and (4) respectively. The details of PL-ADMM are clarified in Section 3.1 and the PL-ADMM-NF algorithm is presented in Section 3.2.
3.1 An Algorithm for SLRQA
@@
1:A noisy observed image $\mathbf{\dot{Y}}$, an initial iterate ${\mathbf{\dot{X}}}_{0}:=\mathbf{\dot{Y}}$, two parameters $\mu\in(0,2)$ and $\beta>0$, a tolerance $\eta>0$, two sequences of regularization parameters $\{L_{k,1}\}$ and $\{L_{k,2}\}$ satisfying $\sup_{k\geq 0}{L_{i,k}}<\infty$ and $\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$.
2:A recovered image ${\mathbf{\dot{X}}}_{*}$.
@@
${\mathbf{\dot{X}}}_{k+1}\in$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\;\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}_{k}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2};$
@@
$\mathbf{\dot{W}}_{k+1}=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}$ $\;\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{W}}-(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+$ $\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}$
@@
In Step 5, ${\mathbf{\dot{X}}}_{k+1}$ is obtained by minimizing $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed $\mathbf{\dot{W}}_{k}$ and $\mathbf{\dot{\Lambda}}_{k}$. This subproblem does not yield a closed-form solution in general and therefore may be solved by nonsmooth algorithms such as a subgradient method or the linearized Bregman algorithm in [6]. If the linear operator $\mathcal{A}$ in (2) is the identity and all pixels of the noisy image $\mathbf{\dot{Y}}$ are observed, i.e., $\Omega$ indicates the whole image, then the ${\mathbf{\dot{X}}}$-subproblem can be solved efficiently. Specifically, the subproblem can be reformulated as
${\mathbf{\dot{X}}}_{k+1}=$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+$ (13)
$\frac{\tau^{2}(\beta+L_{k,1})+1}{2\tau^{2}}\left\|{\mathbf{\dot{X}}}-\frac{\tau^{2}}{\tau^{2}(\beta+L_{k,1})+1}\left(\frac{1}{\tau^{2}}\mathbf{\dot{Y}}+\beta\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k}-\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})+L_{k,1}{\mathbf{\dot{X}}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
If $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sigma_{i}({\mathbf{\dot{X}}}),$ the ${\mathbf{\dot{X}}}$-subproblem has a closed-form solution by using the soft thresholding method on the singular values [6]. When a nonconvex surrogate function $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ is used, the subproblem of ${\mathbf{\dot{X}}}$ can be solved via a convex-concave procedure (CCP) in [22]. For completeness, we summarize CCP in Algorithm 2. Lemma 3.1 is crucial since it transforms the matrix optimization problem of ${\mathbf{\dot{X}}}$ into an optimization problem of singular values. We note that the case of $\varepsilon=0$ has been proven in [12, Theorem 3]. We extend the theorem into the case that $\varepsilon\geq 0$.
Lemma 3.1.
@@
$\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\,\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}$ (14)
@@
$\bm{\sigma_{*}}=\mathop{\mathrm{arg\,min}}_{\bm{\sigma}\,\geq 0}\varphi(\bm{\sigma},\varepsilon;\gamma)+\frac{\mu}{2}\|\bm{\sigma}-\bm{\sigma_{\bm{\hat{{\mathbf{\dot{X}}}}}}}\|_{2}^{2},$ (15)
where $\varphi(\bm{\sigma},\varepsilon;\gamma):=\sum_{i}\phi(\sqrt{\bm{\sigma}_{i}^{2}+\varepsilon^{2}},\gamma)$.
Proof.
According to the definition of QSVD, the following inequality holds.
$\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)=\frac{\mu}{2}\|\bm{\Sigma}-\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ (16)
where the equality is based on the fact that the Frobenius norm is unitarily variant [12, Equation (37)]. Suppose $\mathbf{\dot{A}}=\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}$, thus (16) is equal to
$=\frac{\mu}{2}\|\bm{\Sigma}-\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{A}}}^{*}\mathbf{\dot{A}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ (17)
$\geq\frac{\mu}{2}\left\|\bm{\Sigma}-\bm{\Sigma}_{\mathbf{\dot{A}}}\right\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{A}}}^{*}\mathbf{\dot{A}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
$=\frac{\mu}{2}\left\|\bm{\Sigma}-\bm{\Sigma}_{{\mathbf{\dot{X}}}}\right\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
$=\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}({\mathbf{\dot{X}}})\right\|_{2}^{2}+\varphi(\bm{\sigma}({\mathbf{\dot{X}}}),\varepsilon;\gamma)$
$\geq\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}_{*}\right\|_{2}^{2}+\varphi(\bm{\sigma}_{*},\varepsilon;\gamma),$
where $\bm{\sigma}({\mathbf{\dot{X}}})$ is the singular value vectors of ${\mathbf{\dot{X}}}$, the first inequality follows from Hoffman-Wielandt inequality. According to the definition of ${\mathbf{\dot{X}}}_{*}$, $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}_{*}}^{*}{\mathbf{\dot{X}}}_{*}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}_{*}\|_{\mathrm{F}}^{2}=\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}_{*}\right\|_{2}^{2}+\varphi(\bm{\sigma}_{*},\varepsilon;\gamma)$. The proof is completed. ∎
In the $k$-th iteration of Algorithm 2, the function $\varphi(\bm{\sigma},\varepsilon;\gamma)$ is approximated by a linearized function $\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)+\nabla_{\bm{\sigma}}\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)^{\mathrm{T}}\bm{\sigma}$. It follows that the update of $\bm{\sigma}$ in Problem (15) is given by
$\bm{\sigma}_{k+1}=\max\left\{\bm{\sigma}_{\bm{\hat{{\mathbf{\dot{X}}}}}}-\frac{\nabla_{\bm{\sigma}}\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)}{\mu},\bm{0}\right\}.$ (18)
It has been shown in [22, Theorem 4] that any limit points of $\{\bm{\sigma}_{k}\}_{k=1}^{\infty}$ are stationary points of (15).